# Hypergraph clustering using Ricci curvature: an edge transport perspective

### v1

@@ 4.2 Real data
We evaluate our method on the datasets presented in Table 1. Data and code to reproduce the experiments are available at https://github.com/OlympioH/Ricci_curv_hyp. We adopt the same benchmark study as in Lee and Shin (2023). The method has been evaluated on the following datasets:
  • Co-citation datasets Cora-C, Citeseer and Pubmed where nodes represent research papers, edgess paper co-citing these papers and the communities correspond to similar research thematic.
For each dataset, we iterate Ricci flows $N=20$ times. The edge-cutting threshold $\tau$ is taken as the one that maximizes the hypergraph modularity, computed with the hypernetx python package. We did not observe any significant change of accuracy regarding the weights of the clique and line expansions, as well as in the value of $\alpha$ in the measures defined in Equations 1 and 3. Pairwise curvatures are aggregated at the hyperedge level using the maximum function.
@@
We report the NMI of our method in Table 2, where we compare ourselves with graph and hypergraph partitioning state-of-the-art methods. Node2vec (Grover and Leskovec, 2016), DGI (Veličković et al., 2018) and GRACE (Zhu et al., 2020) embed the hypergraph’s clique expansion in a Euclidean space using various neural networks architectures. $S^{2}$-HHGR from Zhang et al. (2021) and TriCL from Lee and Shin (2023) directly embed the hypergraphs using a neural network architecture. Communities are then detected using a $k$-means algorithm. Note that for TriCL and $S^{2}$-HHGR, additional nodes-feature information is used to improve the prediction, while Ricci flow methods relies only on the hypergraph structural information. All scores for these methods are reported from Lee and Shin (2023). We also report the score obtained by the modularity maximization algorithm from Kamiński et al. (2021). OOM indicates an out-of-memory error on a 24GB GPU (reported from Lee and Shin (2023)). OOT indicates that the results could not be obtained within 72 hours on a single 8-core laptop with a Intel(R) Core(TM) i5-8300H CPU @ 2.30GHz processor unit.
Dataset Cora-C Cora-A Citeseer Pubmed Zoo Mushroom NTU2012
$\sharp$ Nodes 1434 2388 1458 3840 101 8124 2012
$\sharp$ Hyperedges 1579 1072 1079 7963 43 298 2012
Avg. hyperedge size 3.0 4.3 3.2 4.4 39.9 136.3 5.0
Avg. node degree 3.3 1.9 2.4 9.0 17.0 5.0 5.0
$\sharp$ Communities 7 7 6 3 7 2 67
Table 1: Basic statistics of hypergraph data used in the experiments.
@@
N-Ricci 45.8 39.4 38.8 27.8 96.2 OOT 82.3
E-Ricci 43.3 39.1 38.3 22.7 $\mathbf{100.0}$ 42.3 80.9
@@
Ricci-based clustering methods provide an overall competitive accuracy, with a better performance than simple graph vectorization methods, while managing to be as competitive as state-of-the-art neural network-based hypergraph embeddings on a few datasets. In particular, we achieve a perfect clustering on the Zoo dataset using edge-Ricci flow. This dataset stands out for having very few nodes and edges and demonstrates the ability of Ricci-based methods to capture fine structural information on small graphs. We claim that one of the strengths of our method lies in its high level of explainability as opposed to other methods based on neural network embeddings. In addition, methods with the best performance use additional feature information on the nodes while we only use the node-edge incidence information.
4.3 Computational comparison
@@
We observe in Table 2 that nodes-Ricci flow failed to provide results for the Mushroom dataset in a reasonable time. Even after 72 hours, not a single Ricci-flow iteration was completed, despite this dataset containing a relatively small number of hyperedges (see Table 1). To better understand this limitation, we further investigate the computational complexity of both methods.
@@
Computing the Ricci curvature of all edges in a hypergraph $H=(V,E)$ requires solving $|\mathcal{E}|$ optimal transport problems where $\mathcal{E}$ is the edge set of the clique expansion $\mathbf{C}(H)$. In practice, the optimal transport costs between two measures with at most $n$ points are computed using the Sinkhorn algorithm from Sinkhorn (1974), which provides an $\varepsilon$-approximation at a cost $O(n^{2}/\varepsilon^{2})$ according to Lin et al. (2019). This implies worst-case complexities of order $O\left(|\mathcal{E}|\times\underset{x\in V}{\max}|\mathcal{N}(x)|^{2}/\varepsilon^{2}\right)$ for node-Ricci and $O\left(|\mathcal{E}|\times\underset{x\in V}{\max}|St(x)|^{2}/\varepsilon^{2}\right)$ for edge-Ricci. In turn, large hyperedges have a stronger impact on the node-Ricci complexity.
@@
To justify these theoretical claims, we consider a $K$-uniform hypergraph with 1000 nodes and 300 edges of length $K$ drawn uniformly at random. We report the computational times in seconds for a single Ricci-flow interation in Figure 6. Computations are done for unweighted clique and line graphs and averaged over five different hypergraph initializations. We observe that the computational cost of nodes-Ricci curvature indeed blows up for hypergraphs with large hyperedges due to the inflated neighborhood size $|\mathcal{N}(x)|$. Edge-Ricci flow demonstrates much slower growth in computational time, remaining efficient even for hypergraphs with large hyperedges. Computations are executed on a single 8-core laptop with a Intel(R) Core(TM) i5-8300H CPU @ 2.30GHz processor unit. These computations explain why only edges-Ricci could produce results in a reasonable time on the Mushroom dataset. Although this dataset has relatively few hyperedges, the hyperedges are exceptionally large, inducing severe computational challenges for node-Ricci flow. The computational times from Figure 6 along with the quantitative results from Table 2 indicate that Ricci-flow based methods are in particular more suited with dealing with rather small hypergraphs.
@@
Figure 6: Computation times of Ricci curvature for both methods on a $K$-uniform hypergraph as a function of $K$.
@@ 4.4 Wrap-up, comparison of Node-Ricci and Edge-Ricci flows
Based on the findings from Sections 3.3 and 4, node-Ricci and edge-Ricci flows are complementary approaches for competitive hypergraph clustering. Indeed, as the first one primarily leverages the clique expansion and the second one the line expansion information, taken together, they capture most of the information of the hypergraph. More precisely, edges-Ricci is preferable to nodes-Ricci whenever:
  • There are many small communities as discussed in Section 3.3.
  • The intra-community edges are typically smaller than inter-community edges according to Section 4.1.
  • The hypergraph has very large hyperedges, inducing prohibitive computations on the clique graph, see Figure 4.3.
@@ Conclusion
We have developed two methods to extend Ricci flow to hypergraphs. The first one, node-Ricci flow, applies standard Ricci flow on the clique expansion and further aggregates it to hyperedges. The other one, the edge-Ricci flow is the main contribution of this article and constitutes an original way to transport edges using the line expansion. Both methods define new weights on the hyperedges, that can then simply be turned into a partitioning algorithm. Each method has its own advantages depending on the specific characteristics of the hypergraph. A first natural extension of this work would be to consider a co-optimal transport of both nodes and edges, see Titouan et al. (2020); Chowdhury et al. (2024). The notion of discrete Ricci flow has many more practical uses than graph partitioning, see for instance Khodaei et al. (2024). In particular, it can be used as a method to preprocess hypergraphs by generating new weights that better highlight underlying community structures.
Acknowledgments
The author is grateful to Yusuke Imoto, Killian Meehan and Nicolò Zava for valuable discussions. This work has been conducted under a fellowship awarded by the Japanese Society for the Promotion of Science.
@@ Appendix A Proof of Proposition 1
  • We start by computing $w_{\mathcal{N},3}^{(l)}$. Let $x$ and $y$ be two non-gateway nodes from the same clique $(x,y,u_{1},\ldots,u_{a-2})$. We have that $St(x)\setminus(x,y)=\{(x,u_{1}),\ldots,(x,u_{a-2})\}$ and $St(y)\setminus(x,y)=\{(y,u_{1}),\ldots(y,u_{a-2})\}$. The optimal way to transport the uniform probability measure on $St(x)\setminus(x,y)$ to the uniform probability measure on $St(x)\setminus(x,y)$ is to transport each $(x,u_{i})$ to $(y,u_{i})$, which are adjacent in the line graph. Each edge has weight $1/(a-2)$ and there is $a-2$ of them, hence a total cost of $1$. Hence by induction, since $w_{\mathcal{N},3}^{(l)}=1$, we have that $w_{\mathcal{N},3}^{(l)}=1$ for all $l$.
  • For $w_{\mathcal{N},1}^{(l)}$, let $x$ and $y$ be two distinct gateway-nodes, connected by the large hyperedge $E_{b}$. We have $St(x)\setminus E_{b}=\{(x,u_{1}),\ldots,(x,u_{a-1})\}$ and $St(y)\setminus E_{b}=\{(y,v_{1}),\ldots,(x,v_{a-1})\}$. The optimal way to transport the uniform probability measure on $St(x)\setminus E_{b}$ to the uniform probability measure on $St(y)\setminus E_{b}$ is to transport each $(x,u_{i})$ to a corresponding $(y,v_{i})$ via $E_{b}$. This corresponds to a path of length $2$ on the line graph. Each edge has weight $1/(a-1)$ and there is $a-1$ of them, hence a total cost of $2$. As the cost is the same for each pair of nodes in $E_{b}$, the flow of $E_{b}$ is equal to that of any pair $(x,y)\in E_{b}$. We therefore have that $w_{\mathcal{N},3}^{(l)}=2$ for all $l\geq 1$.
  • For $w_{\mathcal{N},2}^{(l)}$, let $x$ be a gateway node and $y$ and non-gateway node from the same clique $(x,y,u_{1},\ldots,u_{a-2}$. We have $St(x)\setminus(x,y)=\{E_{b},(x,u_{1}),\ldots,(x,u_{a-2})\}$ and $St(y)\setminus(x,y)=\{(y,u_{1}),\ldots,(y,u_{a-2})\}$. We represent the line graph restricted to $St(x)\cup St(y)\setminus(x,y)$ in Figure 7. We write the mass of each edge according to Equation 3 where we write $d_{i}=w_{\mathcal{N},i}^{(l-1)}$ to simplify the notation. We have just demonstrated that $d_{1}=2$ and $d_{3}=1$. The optimal transport plan between $\mu_{x}$ and $\mu_{y}$ transports the mass on each $(x,u_{i})$ to $(y,u_{i})$. These edges are adjacent in the line graph. These $(a-2)$ edges therefore occupy a fraction of the total cost of $\frac{(a-2)d_{2}}{2+(a-2)d_{2}}$. The mass $\frac{d_{2}}{2+(a-2)d_{2}}$ on $E_{b}$ is equally transported to each $(y,u_{i})$ through the corresponding $(x,u_{i})$. Each $(y,u_{i})$ is at a distance $2$ from $E_{b}$, see Figure 7. This implies a cost of $\frac{2d_{2}}{2+(a-2)d_{2}}$ for transporting the mass at $E_{b}$. In the end, the total cost of transporting $\mu_{x}$ to $\mu_{y}$ is $\frac{4+(a-2)d_{2}}{2+(a-2)d_{2}}$. This implies that
@@
    Figure 7: Restriction of the line graph to compute the flow between a gateway and a non-gateway node in Proposition 1.

### v2

@@ 4.2 Real data
We evaluate our method on the datasets presented in Table S.1. A detailed description of the datasets can be found in Section B. Data and code to reproduce the experiments are available at https://github.com/AnonRicciHG/Ricci_curv_HG. We adopt the same benchmark study as in Lee and Shin, (2023). For each dataset, we iterate Ricci flows $N=20$ times. The edge-cutting threshold $\tau$ is taken as the one that maximizes the hypergraph modularity, computed with the hypernetx python package.
@@
We report the NMI of our method in Table 1, where we compare ourselves with graph and hypergraph partitioning state-of-the-art methods. Node2vec (Grover and Leskovec,, 2016), DGI (Veličković et al.,, 2018) and GRACE (Zhu et al.,, 2020) embed the hypergraph’s clique expansion in a Euclidean space using various neural networks architectures. $S^{2}$-HHGR from Zhang et al., (2021) and TriCL from Lee and Shin, (2023) directly embed the hypergraphs using a neural network architecture. Communities are then detected using a $k$-means algorithm. Note that for TriCL and $S^{2}$-HHGR, additional node-feature information is used to improve the prediction, while Ricci flow methods rely only on the hypergraph structural information. All scores for these methods are reported from Lee and Shin, (2023). We also report the score obtained by the modularity maximization algorithm from Kamiński et al., (2021). OOM indicates an out-of-memory error on a 24GB GPU (reported from Lee and Shin, (2023)). OOT indicates that the results could not be obtained within 72 hours on a single 8-core laptop with a Intel(R) Core(TM) i5-8300H CPU @ 2.30GHz processor unit.
@@
N-Ricci 45.8 39.4 38.8 27.8 96.2 OOT 76.3
E-Ricci 43.3 39.1 38.3 22.7 $\mathbf{100.0}$ 42.3 77.6
@@
Ricci-based clustering methods provide an overall fair accuracy, with a better performance than simple graph vectorization methods, while managing to be as competitive as state-of-the-art neural network-based hypergraph embeddings on a few datasets. In particular, we achieve a perfect clustering on the Zoo dataset using edge-Ricci flow. This dataset stands out for having very few nodes and edges and demonstrates the ability of Ricci-based methods to capture fine structural information on small graphs. We assert that a significant advantage of our method is its high level of explainability, particularly when compared to methods based on neural network embeddings. Specifically, Ricci flow facilitates clustering on graphs with several notable properties:
  • It is grounded in established theoretical works on metric spaces, see Ollivier, (2007), and benefits from convergence results when applied on random geometric graphs, as shown in Van Der Hoorn et al., (2021).
  • The method is computable in closed-form on simple examples as demonstrated in Ni et al., (2019) and further elaborated in Section $3.$
  • It is strongly associated with the intuitive notion of a community on a graph, by assigning a substantial weight to edges across communities, see Figure 2.
  • The method relies on explicit and transparent computations and is fully deterministic.
However, it is to be noted that when extended to hypergraphs, some of these points fail to be as compelling, notably the theoretical guarantees and the interpretability of the sign of the curvature from Figure 2. Nevertheless, our method’s explicit computation of a flow process, which draws a strong analogy to its well-established geometric counterpart, offers greater interpretability compared to neural network approaches. The latter often involves many parameters, making their interpretation challenging.
4.3 Hyperparameter discussion
Both notions of Ricci curvature on hypergraphs depend on a moderate number of hyperparameters. Some of them can significantly influence the final performance. In this section, we review all the hyperparameters, conduct a sensitivity analysis, and provide guidelines for tuning.
The number of flow iterations NN
This parameter is inherited from Ricci flow on graphs. The effect of this parameter has been examined in Section 4.1 of the Supplementary Material of Ni et al., (2019). The authors experimentally illustrate the convergence of the edge weights as $N$ tend to infinity. Therefore, choosing a large $N$ therefore has no detrimental impact aside from increasing the running time. In the associated code, the authors observe convergence for $N$ between 20 and 50. This guideline still holds in our case, and we did also observe a stabilization of the Ricci flow weights for $N$ larger than 20.
Parameters $\alpha$ and $p$ for $\nu^{\alpha,p}$
These parameters are used to define the measures in Equation 1. These parameters are directly inherited from Ricci flow on graphs. The impact of the parameter $p$ has been studied in Section 4.2 of the Supplementary Material of Ni et al., (2019). We nevertheless decide to review the impact of this parameter in Figure S.2 where we observe that in some cases, there is a drop of performance for $p=0$ while $p=1$ and $p=2$ show a similar performance. Additionally, we decide to investigate the role of the parameter $\alpha$ in some critical cases, see S.3. This study has not been conducted in loc. cit. where the authors advocate for a default choice of $\alpha=0.5$. Although it is challenging to provide specific guidelines, extreme values of the parameter $\alpha$ can lead to serious instabilities. As a result, taking $\alpha$ close to 0.5 is a safe choice.
Aggregation and graph weightings
Both methods compute a Ricci-curvature between pairs of adjacent nodes which is in turn aggregated at the hyperedge level, as suggested in Coupette et al., (2022). We have explored two choices: taking the maximum and the arithmetic average.
In addition, the transport costs depend on the initial clique graph and line graph weightings discussed in Section 3.1. We compare between unweighted graphs and Jaccard weightings. We have studied the joint impact of these two parameters and report the results in Figures S.4 and S.5. We make the following observation:
  • For node-Ricci flow, the clique graph weighting has little to no impact, while using the maximum of the pairwise curvature for the hyperedge curvature consistently yields better results.
  • For edge-Ricci flow, we observe a very important difference between these parameters. In short, the combination Jaccard-maximum is the most competitive when $s_{in}\leq s_{out}$ and Uniform-average is the most competitive when $s_{in}>s_{out}$. The other two combinations always yield an intermediate accuracy. This implies that an intersection-sensitive line graph weighting has a strong positive impact in the case of large hyperedges between communities. Nevertheless, we observe in Figure S.6 that edge-Ricci flow with any combination of parameters is always outperformed by a node-Ricci flow with a maximum aggregation function in the case where $s_{in}>s_{out}$, as already observed in Section 4.1. Therefore, we always recommend using the Jaccard-maximum parameter combination for the edge-Ricci flow to fully leverage its potential in the case of large inter-community hyperedges.
Joint study of every hyperparameter on a real dataset
In the previous paragraphs, we considered critical cases where a change in parameters could have a dramatic impact. We now study the robustness of the two types of transport applied to a real dataset. In Tables S.2 and S.3, we examine the joint effect of every parameter except $\alpha$ and $\tau$ on the dataset NTU2012. We observe that for this real dataset, the choice of weighting has almost no impact on the final accuracy. However, using the maximum function for $Agg$ provides a much better accuracy than the average. Finally, We observe a modest improvement by considering $p=1$ over $p=0$.
Threshold parameter $\tau$
Both notions of Ricci flow simply operates a community-sensitive reweighting of the edges, equivalent to a hierarchical clustering of the nodes. The cut-off parameter $\tau$ enables access to a particular clustering, making it a crucial parameter when comparing to ground-truth labels. To chose the value of $\tau$, we compare the one that maximizes the hypergraph modularity $\tau_{H}$, the one that maximizes the graph modularity on the clique expansion $\tau_{C}$ and the one giving the clustering with the highest NMI, i.e. closest to the ground-truth one $\tau^{\star}$. We report these results in Tables S.4 and S.5. We observe that using $\tau_{H}$ gives results relatively close from the optimal ones using $\tau^{\star}$ and better accuracy than using $\tau_{C}$. In particular, on the Zoo dataset, $\tau_{H}=\tau^{\star}$ for both notions of transport, while $\tau_{C}$ gives extremely poor results. This additional experiment aligns with the findings of Kamiński et al., (2021) and Lee and Shin, (2023), advocating for using hypergraph-specific tools when available, rather than reducing to the clique expansion.
Finally, note that there is not necessarily a single best value for the parameter $\tau$. Indeed, modifying the optimal $\tau$ can provide clusters coarser or finer than the ground-truth, which can be appealing in some applications.
4.4 Computational aspects
Comparison between the two types of transport
@@ 4.4 Computational aspects
We observe in Table 1 that node-Ricci flow failed to provide results for the Mushroom dataset in a reasonable time. Even after 72 hours, not a single Ricci-flow iteration was completed, despite this dataset containing a relatively small number of hyperedges (see Table S.1). To better understand this limitation, we further investigate the computational complexity of both methods.
@@
Computing the Ricci curvature of all edges in a hypergraph $H=(V,E)$ requires solving $|\mathcal{E}|$ optimal transport problems where $\mathcal{E}$ is the edge set of the clique expansion $\mathbf{C}(H)$. In practice, the optimal transport costs between two measures with at most $n$ points are computed using the Sinkhorn algorithm from Sinkhorn, (1974), which provides an $\varepsilon$-approximation at a cost $O(n^{2}/\varepsilon^{2})$ according to Lin et al., (2019). This implies worst-case complexities of order $O\left(|\mathcal{E}|\times\underset{x\in V}{\max}|\mathcal{N}(x)|^{2}/\varepsilon^{2}\right)$ for node-Ricci and $O\left(|\mathcal{E}|\times\underset{x\in V}{\max}|St(x)|^{2}/\varepsilon^{2}\right)$ for edge-Ricci. In turn, large hyperedges have a stronger impact on the node-Ricci complexity.
@@
To justify these theoretical claims, we consider a $K$-uniform hypergraph with 1000 nodes and 300 edges of length $K$ drawn uniformly at random. We report the computational times in seconds for a single Ricci-flow interation in Figure 6. Computations are carried out for unweighted clique and line expansions and are averaged over five different hypergraph initializations. We observe that the computational cost of nodes-Ricci curvature indeed blows up for hypergraphs with large hyperedges due to the inflated neighborhood size $|\mathcal{N}(x)|$. Edge-Ricci flow demonstrates much slower growth in computational time, remaining efficient even for hypergraphs with large hyperedges. Computations are executed on a single 8-core laptop with a Intel(R) Core(TM) i5-8300H CPU @ 2.30GHz processor unit. These computations explain why only edge-Ricci could produce results in a reasonable time on the Mushroom dataset. Although this dataset has relatively few hyperedges, the hyperedges are exceptionally large, inducing severe computational challenges for node-Ricci flow.
@@
Figure 6: Computation times of Ricci curvature for both methods on a $K$-uniform hypergraph with 1000 nodes and 300 edges as a function of $K$.
Comparison on real data
We now compare the two notions of Ricci flow to the modularity-based clustering. As they involve solving many optimal transport problems, flow-based methods appear to be quite costly in practice. We report computational times for $N=20$ flow iterations in Table 2. For datasets which typically have nodes with large stars (especially citation datasets), the computational performance of edge-Ricci can be disadvantageous in some cases when compared to modularity or node-Ricci. The times from Table 2 together with the quantitative results from Table 1 indicate that Ricci-flow based methods are more suited to dealing with rather small hypergraphs to extract fine information on them.
Dataset Cora-C Pubmed NTU2012 Zoo
Modularity 150.1 298 40.1 0.9
N-Ricci 43 7890 48.8 7.4
E-Ricci 296 31500 40.5 17.7
Table 2: Computational time on real data (in seconds)
@@ 4.5 Wrap-up, comparison of Node-Ricci and Edge-Ricci flows
Based on the findings from Sections 3.3 and 4, node-Ricci and edge-Ricci flows are complementary approaches for competitive hypergraph clustering. Indeed, as the first one primarily leverages the clique expansion and the second one the line expansion information, taken together, they capture most of the information of the hypergraph. More precisely, edge-Ricci is preferable to node-Ricci whenever:
  • There are many small communities.
  • The intra-community edges are typically smaller than inter-community edges.
  • The hypergraph has very large hyperedges, inducing prohibitive computations on the clique graph.
@@ Conclusion
We have developed two methods to extend Ricci flow to hypergraphs. The first method, node-Ricci flow, applies standard Ricci flow on the clique expansion and further aggregates it to hyperedges. The second method, edge-Ricci flow, is the main contribution of this article and constitutes an original approach to transporting edges using the line expansion. Both methods define new weights on the hyperedges, which can then be simply turned into a partitioning algorithm. Each method has its own advantages depending on the specific characteristics of the hypergraph.
A natural extension of this work would be to consider a co-optimal transport of both nodes and edges, as discussed in Titouan et al., (2020); Chowdhury et al., (2024). Additionally, the notion of discrete Ricci flow has more practical uses beyond graph partitioning. For instance, it can be used as a method to preprocess hypergraphs by generating new weights that better highlight underlying community structures, as seen in Khodaei et al., (2024). Finally, several notions of curvature have been developed for graph, as explored in Iyer et al., (2024); Cushing et al., (2022). Further extending these notions to hypergraph data constitutes a potential research direction.
@@ Appendix A Proof of Proposition 1
  • We start by computing $w_{\mathcal{N},3}^{(l)}$. Let $x$ and $y$ be two non-gateway nodes from the same clique $(x,y,u_{1},\ldots,u_{a-2})$. We have that $St(x)\setminus(x,y)=\{(x,u_{1}),\ldots,(x,u_{a-2})\}$ and $St(y)\setminus(x,y)=\{(y,u_{1}),\ldots(y,u_{a-2})\}$. Each $(x,u_{i})$ is adjacent to $(y,u_{i})$ and at a distance two from the others $(y,u_{j})_{j\neq i}$? The optimal way to transport the uniform probability measure on $St(x)\setminus(x,y)$ to the uniform probability measure on $St(y)\setminus(x,y)$ is therefore to transport each $(x,u_{i})$ to $(y,u_{i})$, which are at a distance 1 from each other. Each edge has weight 1/(a2)1/(a-2) and there is a2a-2 of them, hence a total cost of $1$. Hence by induction, since $w_{\mathcal{N},3}^{(l)}=1$, we have that $w_{\mathcal{N},3}^{(l)}=1$ for all $l$.
  • For $w_{\mathcal{N},1}^{(l)}$, let $x$ and $y$ be two distinct gateway-nodes, connected by the large hyperedge $E_{b}$. We have $St(x)\setminus E_{b}=\{(x,u_{1}),\ldots,(x,u_{a-1})\}$ and $St(y)\setminus E_{b}=\{(y,v_{1}),\ldots,(x,v_{a-1})\}$. Each $(x,u_{i})$ is at distance at least 2 from a $(y,v_{i})$. Therefore, an optimal way to transport the uniform probability measure on $St(x)\setminus E_{b}$ to the uniform probability measure on $St(y)\setminus E_{b}$ is to transport each $(x,u_{i})$ to a corresponding $(y,v_{i})$ via $E_{b}$, such that the distance is exactly 2. This corresponds to a path of length $2$ on the line graph. Each edge has weight 1/(a1)1/(a-1) and there is a1a-1 of them, hence a total cost of $2$. As the cost is the same for each pair of nodes in $E_{b}$, the flow of $E_{b}$ is equal to that of any pair $(x,y)\in E_{b}$. We therefore have that $w_{\mathcal{N},3}^{(l)}=2$ for all $l\geq 1$.
  • For $w_{\mathcal{N},2}^{(l)}$, let $x$ be a gateway node and $y$ and non-gateway node from the same clique $(x,y,u_{1},\ldots,u_{a-2}$. We have $St(x)\setminus(x,y)=\{E_{b},(x,u_{1}),\ldots,(x,u_{a-2})\}$ and $St(y)\setminus(x,y)=\{(y,u_{1}),\ldots,(y,u_{a-2})\}$. We represent the line graph restricted to $St(x)\cup St(y)\setminus(x,y)$ in Figure S.1. We write the mass of each edge according to Equation 3 where we write $d_{i}=w_{\mathcal{N},i}^{(l-1)}$ to simplify the notation. We have just demonstrated that $d_{1}=2$ and $d_{3}=1$. From Figure S.1, it is clear than an optimal transport plan between $\mu_{x}$ and $\mu_{y}$ transports the mass on each $(x,u_{i})$ to $(y,u_{i})$. These edges are adjacent in the line graph. These (a2)(a-2) edges therefore occupy a fraction of the total cost of $\frac{(a-2)d_{2}}{2+(a-2)d_{2}}$. The mass $\frac{d_{2}}{2+(a-2)d_{2}}$ on $E_{b}$ is equally transported to each $(y,u_{i})$ through the corresponding $(x,u_{i})$. Each $(y,u_{i})$ is at a distance $2$ from $E_{b}$, see Figure S.1. This implies a cost of $\frac{2d_{2}}{2+(a-2)d_{2}}$ for transporting the mass at $E_{b}$. In the end, the total cost of transporting $\mu_{x}$ to $\mu_{y}$ is $\frac{4+(a-2)d_{2}}{2+(a-2)d_{2}}$. This implies that
@@
    Figure S.1: Restriction of the line graph to compute the flow between a gateway and a non-gateway node in Proposition 1.
Appendix B Description of the datasets
The experiments from Section 4.2 have been conducted on these hypergraph datasets:
  • Co-citation datasets Cora-C, Citeseer and Pubmed where nodes represent research papers, edges papers co-citing these papers and the communities correspond to similar research thematic.
We refer to Table S.1 for a quantitative description of each dataset.
Dataset Cora-C Cora-A Citeseer Pubmed Zoo Mushroom NTU2012
$\sharp$ Nodes 1434 2388 1458 3840 101 8124 2012
$\sharp$ Hyperedges 1579 1072 1079 7963 43 298 2012
Avg. hyperedge size 3.0 4.3 3.2 4.4 39.9 136.3 5.0
Avg. node degree 3.3 1.9 2.4 9.0 17.0 5.0 5.0
$\sharp$ Communities 7 7 6 3 7 2 67
Table S.1: Basic statistics of hypergraph data used in the experiments.
Appendix C Impact of the hyperparameters
Unless specified otherwise, we report the NMI in the synthetic set-up of Section 4.1. Results are averaged over 5 random hypergraph generation and comparison with the ground-truth is realized for the optimal cut-off parameter $\tau^{\star}$.
Measure parameter $p$.
In Figure S.2, we report the NMI for three values $p=0,1,2$ for the nodes transport, compared with the edge transport. We set $Agg$ to be the maximum function and consider a uniform weighting of the clique graph. We can see that in some cases, there is a clear drop of accuracy for $p=0$, while $p=1$ and $p=2$ display similar performances.
(a) $s_{in}=2,s_{out}=3$
(b) $s_{in}=3,s_{out}=3$
(c) $s_{in}=4,s_{out}=3$
(d) $s_{in}=2,s_{out}=5$
(e) $s_{in}=3,s_{out}=5$
(f) $s_{in}=4,s_{out}=5$
Figure S.2: Study on the impact of the parameter $p$ to cluster hypergraph stochastic block models.
Measure parameter $\alpha$.
In Figure S.3, we report the NMI for a few specific critical values of $s_{in},s_{out},N_{in},N_{out}$. These results are obtained for $p=1$ for the node transport. We set $Agg$ to be the maximum function and consider a uniform weighting of the clique graph. We can see that in these critical cases, the parameter $\alpha$ can play a very big impact. It is difficult to provide any generic guideline, but extreme values of $\alpha$ seem to provoke instabilities in some cases.
(a) $s_{in}=3,s_{out}=6$
(b) $s_{in}=3,s_{out}=4$
(c) $s_{in}=3,s_{out}=4$
Figure S.3: Study on the impact of the parameter $\alpha$ to cluster hypergraph stochastic block models
Aggregation and clique graph weighting
In Figure S.4 we report the NMI in the synthetic set-up of Section 4.1 where we simultaneously compare the effect of the clique graph weighting and the curvature aggregation at the hyperedge level. We compare the four possible combinations, where the clique graph can have either a Jaccard or a uniform weighting, and $Agg$ can be either the maximum or the arithmetic average. We do not notice any significant difference between the two types of clique graph weightings. However, we observe that using the maximum pairwise curvature instead of the average yields similar to much better results and should therefore always be preferred.
(a) $s_{in}=2,s_{out}=5$
(b) $s_{in}=3,s_{out}=5$
(c) $s_{in}=5,s_{out}=4$
Figure S.4: Study on the impact of the clique graph weighting and the aggregation function to cluster hypergraph stochastic block models with node-Ricci flow
Aggregation and line graph weighting
In Figure S.5 we report the NMI in the synthetic set-up of Section 4.1 where we simultaneously compare the effect of the line graph weighting and the curvature aggregation at the hyperedge level. We compare the four possible combinations, where the line graph can have either a Jaccard or a uniform weighting, and $Agg$ can be either the maximum or the arithmetic average. The differences can be very significative, in particular between the combinations Jaccard-maximum and Uniform-average. Overall, we observe that the former provides a much better clustering accuracy whenever $s_{in}\leq s_{out}$ while the latter is better when $s_{in}>s_{out}$. However, we observe in Figure S.6 that in this case, edge-Ricci flow with any parameter still underperforms compared to node-Ricci flow, which is particularly adapted to dealing with the situation of large intra-community hyperedges and small inter-community hyperedges. We consider a node-Ricci flow with parameters Jaccard-maximum. As a result, we claim that the parameter combination Uniform-average for edge-flow should never be preferred as it will always be beaten by either the Jaccard-maximum edge-flow or by the node-Ricci flow.
(a) $s_{in}=2,s_{out}=4$
(b) $s_{in}=4,s_{out}=2$
(c) $s_{in}=4,s_{out}=4$
Figure S.5: Study on the impact of the line graph weighting and the aggregation function to cluster hypergraph stochastic block models with edge-Ricci flow
(a) $s_{in}=2,s_{out}=4$
(b) $s_{in}=4,s_{out}=2$
(c) $s_{in}=5,s_{out}=3$
Figure S.6: Comparison of Jaccard-maximum and Uniform-average edge-Ricci flow with node-Ricci flow.
Joint study of every hyperparameter on a real dataset.
We study the impact of all parameters except $\alpha$ and $\tau$ for the clustering task on the dataset NTU2012. The clustering accuracies are reported for the optimal $\tau^{\star}$. Results are reported in Tables S.2 and S.3. We can see that in this real data case, The choice of the weights for the clique graph and the line graph has a negligible impact. For both types of Ricci flow, considering the maximum curvature over the average one has a tremendous impact. In addition, we can observe an improvement by considering $p=1$ over $p=0$. For $p=2$, numerical instabilities prevented the code from compiling.
Aggregation Clique graph weights $p$ NMI
Max Jaccard 0 79.7
Max Uniform 0 79.7
Average Jaccard 0 69.1
Average Uniform 0 69.1
Max Jaccard 1 82.1
Max Uniform 1 82.5
Average Jaccard 1 69.1
Average Uniform 1 69.1
Table S.2: Node-Ricci flow clustering of the NTU2012 dataset for various sets of hyperparameters
Aggregation Clique graph weights NMI
Max Jaccard 80.2
Max Uniform 80.2
Average Jaccard 69.1
Average Uniform 69.1
Table S.3: Edge-Ricci flow clustering of the NTU2012 dataset for various sets of hyperparameters
Threshold parameter $\tau$
We now try to assess the estimation of the $\tau^{\star}$ parameter providing the highest clustering accuracy. We compare the NMI obtained by trimming edges with a flow greater than $\tau^{\star}$ with the NMI obtained with $\tau_{H}$ and $\tau_{C}$. We report the results for the clustering of the datasets NTU2012 and Zoo in Tables S.4 and S.5. Based on the previous findings, we considered a Jaccard weighting, a hyperedge aggregation with the $\max$ function, and $p=1$ for the nodes transport. On the NTU2012 dataset, we observe a better performance using the hypergraph notion of modularity over the clique graph modularity. The loss of accuracy for Edge-Ricci flow is milder than for Node-Ricci flow when using an approximate $\tau$ parameter. For the Zoo dataset, we observe a very strong difference between the two notions of modularity, as considering the hypergraph modularity yields the optimal clustering, while the modularity on the clique graph gives extremely poor results. This has to be paralleled with the results from Table 1 where methods relying only on the clique graph expansion provided very unsatisfactory results.
Type of flow $\tau^{\star}$ $\tau_{H}$ $\tau_{C}$
Nodes 82.1 76.3 72.3
Edges 80.2 77.6 75.2
Table S.4: NMI for various choices of $\tau$ on the NTU2012 dataset
Type of flow $\tau^{\star}$ $\tau_{H}$ $\tau_{C}$
Nodes 96.2 96.2 18.3
Edges 100 100 18.3
Table S.5: NMI for various choices of $\tau$ on the Zoo dataset